# Automated Modeling Method for Pathloss Model Discovery

### v3

@@ VI-B Learning the CI Model from Synthetic Data
Table VI presents the performance comparison of various methods for discovering $PL^{CI}$. The KAN [4,4,1] method achieved the best results, with minimal error values (MAE = 0.05, MSE = 0.01, MAPE = 0.04) and an $R^{2}$ value of 1, showing a highly accurate fit between the predicted values and the actual values as illustrated in Figure 6. In the figure, pred represents the predicted PL values and true represents the actual values based on Eq. (2). The KAN [4,4,1] Auto-symbolic showed lower performance, with MAE = 4.56, MSE = 36.26, MAPE = 3.05, and $R^{2}$ = 0.96. Among the DSR methods, DSR-RSPG yielded the lowest error (MAE = 11.49, MSE = 261.46, MAPE = 10.21, $R^{2}$ = 0.80), followed by DSR-PQT (MAE = 13.60, $R^{2}$ = 0.84) and DSR-VPG (MAE = 16.90, $R^{2}$ = 0.67). These results demonstrate that the KAN [4,4,1] significantly outperforms the other methods, with smaller errors corresponding to better predictive accuracy.
@@
TABLE VI: Performance comparison on CI model.
Methods Metrics MAE MSE MAPE R2 Expression Interpretability Validity
KAN [4,4,1]* 0.05 0.01 0.04 1 Tree and splines in Figure 7 Medium ?
KAN [4,4,1] Auto-symbolic* 4.56 36.26 3.05 0.96 $-280.5\operatorname{log}_{10}{\left(8.7-1.6n\right)}+43.8\operatorname{log}_{10}{\left(1.1d+1.9\right)}+19.9\operatorname{log}_{10}{\left(5.2f+9.3\right)}+50.8\operatorname{log}_{10}{\left(0.8\chi_{\sigma}+7.5\right)}+352.2$ High
DSR-RSPG 11.49 261.46 10.21 0.80 $23n+\frac{d}{10}+log_{10}(f)+40$ High
DSR-PQT 13.60 223.05 10.58 0.84 $23n+\frac{d}{10}+(\chi_{\sigma}+n)cos(log_{10}(f))+40$ High
DSR-VPG 16.90 570.92 15.24 0.67 $21n+2\chi_{\sigma}+60$ High
@@
Figure 7 shows the KAN architecture utilized for $PL^{CI}$ modeling, featuring a three-layer design [4, 4, 1]. The first layer consists of four input nodes corresponding to the parameters $f,n,d,\chi_{\sigma}$, followed by a hidden layer with four nodes, and a single-node output layer representing the predicted function value.
From the learned graph in Figure 7, it can be seen that the final expression is the sum of three terms, where the second has a lower weight. The rightmost node is disconnected because the model has learned that the connecting edges are not important. The functions associated with the first two edges of the last layer apply nonlinear transformations, whereas the third edge performs a quasi-linear transformation. From the coloring of the edges and summations in the hidden layer, it can be observed that, in the first summation node, a quasi-linear transformation of $n$ and a nonlinear transformation of $d$ are combined, both with medium weighting. The second summation node aggregates the nonlinear transformation of $n$ and $d$ with low weights. Finally, the third summation node aggregates all four inputs with high to medium weights.
@@ VI-C Learning an Indoor Model from Empirical Data
The performance comparison of various methods for modeling PL in an indoor environment is presented in Table VII. The $PL^{MWF}$ serves as the baseline, yielding the highest errors (MAE = 9.01, MSE = 121.54, MAPE = 8.98) and $R^{2}$ value of 0.56, indicating a moderate predictive performance. $PL^{EI}$, based on Eq. (4) from [ElChall], slightly improves accuracy, with MAE = 7.64, MSE = 111.67, MAPE = 7.52, and $R^{2}$ = 0.49. However, all automated methods outperform both the baseline and the empirical model. KAN [4,1] significantly improves accuracy (MAE = 6.11, MSE = 59.87, MAPE = 6.22) and a substantially higher $R^{2}$ value of 0.76, as shown in Figure 8, where pred stands for the predicted PL values and true stands for the actual measurements from [ElChall]. The KAN [4,1] Auto-Symbolic performs similarly, with slightly higher errors but maintaining the same $R^{2}$ value. Among DSR methods, DSR-RSPG achieves the overall superior performance (MAE = 5.95, MSE = 55.73, MAPE = 5.96, $R^{2}$ = 0.71), though its $R^{2}$ remains inferior to that of the KAN models. DSR-PQT and DSR-VPG demonstrate comparatively weaker performance, with higher errors and lower values of $R^{2}$ (0.65 and 0.57, respectively). Overall, the results highlight that both KANs and DSR models provide substantial improvements over the baseline model and the indoor empirical model.
Interpretability
Figure 9 illustrates the architecture of the KAN [4,1] used for indoor PL prediction. The architecture has two layers; the first layer consists of four input nodes corresponding to the parameters $n_{w},n_{f},d,f$, and a single-node output layer representing the predicted value. As illustrated in the figure, the final expression is a sum of four terms, with the fourth term contributing less due to its smaller weight. Notably, the function learned on the first edge applies a nonlinear transformation, whereas the remaining three edges implement quasi-linear transformations.
@@
TABLE VII: Performance comparison in an indoor environment.
Methods Metrics MAE MSE MAPE R2 Expression Interpretability Validity
KAN [4,1] 6.11 59.87 6.22 0.76 learned tree and splines as in Figure 9. Medium ?
KAN [4,1] Auto-symbolic 6.27 61.62 6.38 0.76 $41.5\log_{10}\left(0.4d+4.03\right)-1.3\cos\left(5.04f-0.5\right)$ $4.9\sin\left(8.2n_{w}-2.9\right)+52.03\log_{10}\left(2.4n_{f}+8.6\right)-11.2$ High
DSR-RSPG 5.95 55.73 5.96 0.71 $\frac{d}{2}+3log_{10}(f)+5n_{w}+70$ High
DSR-PQT 7.09 80.99 7.19 0.65 $\frac{d}{2}+4n_{w}+80$ High
DSR-VPG 7.38 84.11 7.29 0.57 $n_{f}^{2}+4n_{f}+n_{f}n_{w}+80$ High
@@
Validity
For the indoor environment, the KAN [4,1] Auto-symbolic expression contains a $\cos(f)$ term, implying a periodic dependence on frequency. This contradicts the physical behavior of the signal, as PL increases consistently with $f$. In contrast, the expression obtained by DSR-RSPG is valid, as it includes both $f$ and $d$. Furthermore, the findings highlight the invalidity of the expressions of PQT and VPG: the expression of DSR-PQT does not include $f$, while the expression of DSR-VPG excludes both $f$ and $d$.
@@ VI-D Learning an Outdoor Model from Empirical Data
Table VIII summarizes the performance of different methods for PL prediction in an outdoor environment. The $PL^{FS}$, yields high error values (MAE = 34.08, MSE = 1242.99, MAPE = 31.43) and a low $R^{2}$ value of 0.30, indicating poor predictability. The $PL^{EO}$, represented by Eq. (5) from [ElChall], substantially improves accuracy (MAE = 18.09, MSE = 398.88, MAPE = 17.82), but retains the same $R^{2}$ value of 0.30. In contrast, the KAN model [3,1] achieves significantly better performance, with the lowest errors (MAE = 4.59, MSE = 38.59, MAPE = 4.36) and a higher $R^{2}$ value of 0.71, as demonstrated in Figure 10 showing the comparison between predicted and measured PL values. The KAN [3,1] Auto-symbolic yields poor performance (MAE = 9.50, $R^{2}$ = 0.01), showing a higher error and a lower fit compared to the original KAN structure. Furthermore, the very low $R^{2}$ score shows low interdependence between the input and output variables, meaning that the function to spline matching process resulted in an inaccurate match, leading to significant degradation relative to the KAN-only network.
Among the DSR models, DSR-RSPG performs the best (MAE = 6.62, MSE = 72.09, MAPE = 6.31); however, its $R^{2}$ value (0.29) remains low. DSR-PQT and DSR-VPG show slightly higher errors and comparable $R^{2}$ values. Overall, the KAN [3,1] model outperforms both traditional and DSR methods, offering a more accurate method for outdoor PL prediction.
@@
Figure 11 illustrates the KAN architecture evaluated for the outdoor model, consisting of a two-layer design [3, 1]. The first layer includes three input nodes corresponding to the parameters $h_{\text{ED}},d,f$, followed by a single output node representing the predicted PL value. The figure shows that the final expression is a sum of three terms, with the first and third terms having smaller weights. It can be observed that the functions learned on the second and third edges apply nonlinear transformations, whereas the first implements a quasi-linear transformation.
@@
TABLE VIII: Performance comparison in an outdoor environment.
Methods Metrics MAE MSE MAPE R2 Expression Interpretability Validity
KAN [3,1] 4.59 38.59 4.36 0.71 learned tree and splines as in Figure 11 Medium ?
KAN [3,1] Auto-symbolic 9.50 133.42 9.34 0.01 $\cos{\left(8.7f-3.5\right)}-2.3\operatorname{log}_{10}{\left(7.7-2.2h_{\text{ED}}\right)}+105.8$ High
DSR-RSPG 6.62 72.09 6.31 0.29 $h_{\text{ED}}d+\log_{10}(f)+80$ High
DSR-PQT 7.25 80.45 6.78 0.29 $\log_{10}(\frac{d}{h_{\text{ED}}})+\frac{d}{10}+90$ High
DSR-VPG 8.18 97.90 7.46 0.28 $\frac{d+sin(f)}{10}+90$ High
* Expressions developed with traditional methods used as baselines.
@@
The only valid expression for the outdoor environment is obtained from DSR-RSPG, as it includes both $f$ and $d$. The KAN [3,1] Auto-symbolic expression excludes $d$ and incorporates a cosine function of $f$, which implies periodicity that is not the case with pathloss models. The lack of validity introduced by this cosine component is also confirmed by the very low $R^{2}$ score. The expression of DSR-PQT excludes $f$. Finally, although the DSR-VPG expression includes both $f$ and $d$, it defines $f$ using a sine function, introducing a periodicity that misrepresents the actual relationship between $f$ and PL.
@@ VI-E Discussion
KANs consistently achieve superior predictive accuracy across all scenarios. For example, KAN [6,6,1] and KAN [4,4,1] achieve exceptionally low error metrics and $R^{2}$ values of 1, emphasizing their superior ability to model complex data relationships. The learned models, in the form of spline approximations and weights in a graph composition, are more interpretable compared to deep learning alternatives; however, they are not conventional human-interpretable expressions. To gain insights into the interpretability of the KANs, auto-symbolic is utilized to map from spline-based structures to more interpretable functional forms.
@@
Auto-symbolic variants, such as KAN [6,6,1] Auto-symbolic and KAN [3,1] Auto-symbolic, exhibit invalid expressions, higher error rates, and, in some cases, notably lower $R^{2}$ values compared to their original forms. This illustrates the limits of the current mapping capabilities of KANs, resulting in less accurate and interpretable expressions. In contrast, DSR methods, especially DSR-RSPG, generate simple, valid, and interpretable expressions. Although DSR methods achieve a reasonable level of accuracy, their predictive performance is generally outperformed by KANs. The lower values observed in the DSR methods indicate a trade-off in which predictive precision is reduced to preserve the simplicity and interpretability of the resulting expressions.
@@
Overall, our findings indicate that KANs deliver highly accurate results with superior first-level interpretability(spline-based structures) compared to typical DNN activation maps. However, when it comes to the second level of interpretability, where actual functions are extracted using the Auto-symbolic, KANs are outperformed by certain DSR methods. On the other hand, DSR methods prioritize interpretability over predictive accuracy. The choice between these methods should be determined by the intended objective, whether to maximize predictive accuracy or to derive interpretable expressions that offer clear analytical insight.

### v4

@@ VI-B Learning the CI Model from Synthetic Data
[…]
@@
[…]
@@
Figure 9 shows the KAN architecture utilized for $PL^{CI}$ modeling, featuring a three-layer design [4, 4, 1]. The first layer consists of four input nodes corresponding to the parameters $f,n,d,\chi_{\sigma}^{\text{CI}}$, followed by a hidden layer with four nodes, and a single-node output layer representing the predicted function value.
From the learned graph in Figure 9, it can be seen that the final expression is the sum of three terms, where the second has a lower weight. The rightmost node is disconnected because the model has learned that the connecting edges are not important. The functions associated with the first two edges of the last layer apply nonlinear transformations, whereas the third edge performs a quasi-linear transformation. From the coloring of the edges and summations in the hidden layer, it can be observed that, in the first summation node, a quasi-linear transformation of $n$ and a nonlinear transformation of $d$ are combined, both with medium weighting. The second summation node aggregates the nonlinear transformation of $n$ and $d$ with low weights. Finally, the third summation node aggregates all four inputs with high to medium weights. Looking at the DL baselines, only TabNet provides a low-level interpretability via feature importance scores (f $\approx$ 0.29, n $\approx$ 0.38, d $\approx$ 0.21, $\chi_{\sigma}^{\text{CI}}\approx$ 0.11) without revealing the explicit functional relationships to the PL.
@@
TABLE VII: Performance comparison in an indoor environment. The interpretability and validity assessment follow the methodology in Section V-D.
[…]
KAN [4,1] 5.89 $\pm$0.13 57.67 $\pm$3.30 5.84 $\pm$0.12 0.77 $\pm$0.02 Learned tree and splines as in Figure 11 Medium ?
KAN [4,1] Auto-symbolic 6.33 $\pm$0.23 63.51 $\pm$5.41 6.33 $\pm$0.21 0.74 $\pm$0.02 $41.5\log_{10}\left(0.4d+4.03\right)-1.3\cos\left(5.04f-0.5\right)$ $4.9\sin\left(8.2n_{w}-2.9\right)+52.03\log_{10}\left(2.4n_{f}+8.6\right)-11.2$ High
DSR-RSPG 6.58 $\pm$0.23 67.85 $\pm$1.90 6.69 $\pm$0.29 0.73 $\pm$0.01 $\frac{d}{2}+3\log_{10}(f)+5n_{w}+70$ High
DSR-PQT 6.48 $\pm$0.10 66.11 $\pm$2.07 6.59 $\pm$0.14 0.74 $\pm$0.01 $\frac{d}{2}+4n_{w}+80$ High
DSR-VPG 6.39 $\pm$0.12 64.25 $\pm$2.39 6.48 $\pm$0.18 0.75 $\pm$0.01 $n_{f}^{2}+4n_{f}+n_{f}n_{w}+80$ High
ResNet-MLP 5.67 $\pm$0.38 54.97 $\pm$7.71 5.54 $\pm$0.36 0.78 $\pm$0.03 / None
TabNet 5.14 $\pm$0.27 44.02 $\pm$5.01 5.09 $\pm$0.28 0.82 $\pm$0.02 Feature importance: $n_{w}$: 0.1332, $n_{f}$: 0.2643, $d$: 0.5505, $f$: 0.0520 Low
* Expressions developed with traditional methods used as baselines.
@@ VI-C Learning an Indoor Model from Empirical Data
[…]
@@
[…]
Interpretability
Figure 11 illustrates the architecture of the KAN [4,1] used for indoor PL prediction. The architecture has two layers; the first layer consists of four input nodes corresponding to the parameters $n_{w},n_{f},d,f$, and a single-node output layer representing the predicted value. As illustrated in the figure, the final expression is a sum of four terms, with the fourth term contributing less due to its smaller weight. Notably, the function learned on the first edge applies a nonlinear transformation, whereas the remaining three edges implement quasi-linear transformations. Despite their superior accuracy, neither DL baseline offers a transparent functional form. ResNet-MLP acts as a black box, while TabNet provides limited interpretability through feature importance ($d\approx 0.55$, $n_{f}\approx 0.26$, $n_{w}\approx 0.13$, $f\approx 0.05$).
Validity
For the indoor environment, the KAN [4,1] Auto-symbolic expression contains a $\cos(f)$ term, implying a periodic dependence on frequency. Physically, PL represents the average signal attenuation caused by wavefront spreading and interactions with the environment [9] (e.g., reflection, diffraction, and material penetration). These phenomena make the attenuation typically increase monotonically with frequency and distance. Therefore, oscillatory forms like cosine or sine functions are characteristics of small-scale fading rather than large-scale fading, i.e., PL discussed in this work, are inappropriate to be used for modeling PL. To prevent such anomalies, physics-guided constraints should be applied, such as limiting the search space to monotonic operators. In contrast, the expression obtained by DSR-RSPG is valid, as it includes both $f$ and $d$. Furthermore, the findings highlight the invalidity of the expressions of PQT and VPG: the expression of DSR-PQT does not include $f$, while the expression of DSR-VPG excludes both $f$ and $d$.
TABLE VIII: Performance comparison in an outdoor environment. The interpretability and validity assessment follow the methodology in Section V-D.
[…]
KAN [3,1] 4.83 $\pm$0.36 50.54 $\pm$12.78 4.66 $\pm$0.37 0.57 $\pm$0.12 Learned tree and splines as in Figure 13 Medium ?
KAN [3,1] Auto-symbolic 7.73 $\pm$0.35 99.86 $\pm$6.05 7.53 $\pm$0.43 0.15 $\pm$0.04 $\cos{(8.7f-3.5)}-2.3\log_{10}{(7.7-2.2h_{\text{ED}})}+105.8$ High
DSR-RSPG 7.69 $\pm$0.24 95.60 $\pm$6.70 7.52 $\pm$0.26 0.28 $\pm$0.05 $h_{\text{ED}}d+\log_{10}(f)+80$ High
DSR-PQT 7.38 $\pm$0.19 89.02 $\pm$5.00 7.20 $\pm$0.21 0.33 $\pm$0.04 $\log_{10}(\frac{d}{h_{\text{ED}}})+\frac{d}{10}+90$ High
DSR-VPG 8.02 $\pm$0.24 101.66 $\pm$5.91 7.87 $\pm$0.28 0.23 $\pm$0.04 $\frac{d+\sin(f)}{10}+90$ High
ResNet-MLP 7.87 $\pm$0.43 101.74 $\pm$9.74 7.48 $\pm$0.44 0.14 $\pm$0.09 / None
TabNet 7.44 $\pm$0.42 90.69 $\pm$8.69 7.21 $\pm$0.44 0.23 $\pm$0.06 Feature importance: $h_{\text{ED}}$: 0.2566, $d$: 0.5952, $f$: 0.1481 Low
@@ VI-D Learning an Outdoor Model from Empirical Data
[…]
Among the DSR models, DSR-PQT performs the best (MAE = 7.38 $\pm$0.19, MSE = 89.02 $\pm$5.00, MAPE = 7.20 $\pm$0.21, $R^{2}$ = 0.33 $\pm$0.04); however, its $R^{2}$ value remains low. DSR-RSPG and DSR-VPG show slightly higher errors and lower $R^{2}$ values.
[…]
@@
Figure 13 illustrates the KAN architecture evaluated for the outdoor model, consisting of a two-layer design [3, 1]. The first layer includes three input nodes corresponding to the parameters $h_{\text{ED}},d,f$, followed by a single output node representing the predicted PL value. The figure shows that the final expression is a sum of three terms, with the first and third terms having smaller weights. It can be observed that the functions learned on the second and third edges apply nonlinear transformations, whereas the first implements a quasi-linear transformation. Regarding DL baselines interpretability, similar to other datasets, ResNet-MLP again functions as a complete black box, whereas TabNet provides limited transparency via feature importance ($d\approx 0.60$, $h_{\text{ED}}\approx 0.26$, $f\approx 0.15$).
@@
The only valid expression for the outdoor environment is obtained from DSR-RSPG, as it includes both $f$ and $d$. The KAN [3,1] Auto-symbolic expression excludes $d$ and incorporates a cosine function of $f$, As noted in the indoor case, this introduces a periodic behavior that violates the physics that PL should show a monotonous decay trend over large distances ($d$) and frequencies ($f$). The expression of DSR-PQT excludes $f$. Finally, although the DSR-VPG expression includes both $f$ and $d$, it defines $f$ using a sine function, introducing a periodicity that misrepresents the actual relationship between $f$ and PL.
@@ VI-E Discussion
KANs consistently achieve superior predictive accuracy across all scenarios. For example, KAN [6,6,1] and KAN [4,4,1] achieve exceptionally low error metrics and $R^{2}$ values of 1, emphasizing their superior ability to model complex data relationships. The learned models, in the form of spline approximations and weights in a graph composition, are more interpretable compared to DL alternatives; however, they are not conventional human-interpretable expressions. To gain insights into the interpretability of the KANs, auto-symbolic is utilized to map from spline-based structures to more interpretable functional forms as illustrated in Figure 2 and discussed in Section V-B.
@@
Auto-symbolic variants, such as KAN [6,6,1] Auto-symbolic and KAN [3,1] Auto-symbolic, exhibit invalid expressions, higher error rates, and, in some cases, notably lower $R^{2}$ values compared to their original forms. This illustrates the limits of the current mapping capabilities of KANs, resulting in less accurate and interpretable expressions. In contrast, DSR methods, especially DSR-RSPG, generate simple, valid, and interpretable expressions. Although DSR methods achieve a reasonable level of accuracy, they are computationally more expensive and their predictive performance is generally outperformed by KANs. The lower values observed in the DSR methods indicate a trade-off in which predictive precision is reduced to preserve the simplicity and interpretability of the resulting expressions.
While in this section, the paper analyzed the performance, functional interpretability, and validity of the proposed AutoPL method, with KANs and DSRs while also adding ResNet-MLP and TabNet as alternative neural network baselines, a comprehensive computational cost vs performance Pareto front study is left for future. As AI tools for scientific discovery, including KAN and DSR-based, essentially explore a large search space to find the right distribution for the phenomenon underlying the observed data, the methods themselves are subject to their own hyperparameter settings. To provide insight into this complexity, we considered $R^{2}$ vs selected parameters for both KANs in Section V-B and DSR in Section V-C.
@@
Overall, our findings indicate that KANs deliver highly accurate results with superior first-level interpretability (spline-based structures) compared to typical DNN activation maps. However, when it comes to the second level of interpretability, where actual functions are extracted using the Auto-symbolic, KANs are outperformed by certain DSR methods. On the other hand, DSR methods can be constrained to automatically prioritize interpretability over predictive accuracy while KANs involve slightly more manual labor to achieve that. In our experiments that can be reproduced using the available open source code, we noticed that DSRs tend to take longer to train compared to KANs. The choice between these methods should be determined by the intended objective, whether to maximize predictive accuracy or to derive interpretable expressions that offer clear analytical insight. Generally, the selection and validation of the simplest and most functionally interpretable expression remains manual, as illustrated in Figure 1, with many potential candidates suggested by AutoPL. Future development to introduce physics constraints during training might speed up also that part.
VI-F Scope and Limits of Interpretability
This work emphasizes explicit functional interpretability rather than physical-mechanistic. AutoPL yields closed-form expressions for PL as a function of distance, frequency, wall/floor counts, and antenna height. These expressions expose parameter sensitivities and monotonic trends, enable direct comparison with established models (e.g., CI and ABG), and can be integrated into system-level analysis and planning tools.
The resulting models are not intended to disentangle specific propagation mechanisms, such as reflection, diffraction, scattering. They should therefore be interpreted as empirical, data-driven surrogates, similar in spirit to CI/ABG, that describe what best fits the observed data rather than providing a first-principles electromagnetic account of why propagation behaves as it does.
This aligns with the philosophy of large-scale PL modeling, which favors compact, analytically tractable expressions over physical granularity. Within this framework, AutoPL automates the discovery of such expressions directly from data without imposing predefined model structures.
VI-G Limitations and Guidelines for Safe Adoption
While AutoPL provides a powerful framework for the automated discovery of PL models, its deployment in real-world network planning requires an understanding of its inherent limitations. To ensure the safe adoption of the discovered models, we identify the following constraints:
1) Dataset Scope and Bias: The accuracy of a discovered symbolic model is strictly bounded by the diversity of the training data. If the input dataset is limited to a specific frequency band, such as sub-6 GHz, or a specific environment (such as Urban Micro), the framework may discover an empirical relationship that captures environment-specific noise rather than a generalizable physical law. Users should ensure that the dataset spans the full operational range of the intended application.
2) Sensitivity of Auto-Symbolic Mapping: The transition from a high-dimensional search space (in DSR) or a flexible activation function (in KAN) to a closed-form symbolic expression involves a trade-off between mathematical simplicity and predictive fidelity. There is a risk of ”over-simplification,” where a discovered model ignores subtle physical inflection points—such as the break-point distance in two-slope models—to achieve a higher ”simplicity score.” We recommend users inspect the Pareto front of complexity versus accuracy provided by AutoPL before selecting a final model.
3) Domain Shift and Generalization: Like most data-driven techniques, models discovered via AutoPL are susceptible to performance degradation when subjected to domain shift. A model discovered in a specific geographical region may fail to generalize to a different city due to variations in building materials, foliage density, or atmospheric conditions. Consequently, a discovered model should be treated as a ”specialized” expression for its target domain unless validated against diverse out-of-distribution (OOD) datasets.
While the current AutoPL framework does not enforce physical constraints during optimization, future extensions could incorporate physics-informed constraints directly into the learning objective. Examples include enforcing monotonicity with respect to distance and frequency and restricting the functional search space to non-oscillatory operators. Such constraints could reduce the occurrence of physically implausible expressions while preserving the flexibility of automated model discovery.
4) Operational Safeguards: To adopt AutoPL safely, we include a ”Human-Scientist-in-the-Loop” validation step as depicted in Figure 1.